rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#1822 — FS#5796 — Optical loop (LIB59 SEC59 FAC59 BOI59 BLE59 PIL59)

Attached to Project— ADSL / SDSL
Maintenance
Backend / Core
CLOSED
100%
We are doing working tonight at midnight in order to replace the optical fibers. An interruption of a few minutes is expected on each of the sites.
We will update regularly this task in order to keep you informed of the task's evolution.
Date:  Tuesday, 20 September 2011, 07:19AM
Reason for closing:  Done
Comment by OVH - Tuesday, 20 September 2011, 00:49AM

We are unplugging PIL59 in order to rewire it. A few minutes cut is expected.


Comment by OVH - Tuesday, 20 September 2011, 00:49AM

PIL59 fiber is reconnected, we reconfigure the routing.


Comment by OVH - Tuesday, 20 September 2011, 00:50AM

The PIL site is UP. We are moving forward with works!


Comment by OVH - Tuesday, 20 September 2011, 00:51AM

The site of LIB59 will be disconnected in a few minutes.


Comment by OVH - Tuesday, 20 September 2011, 02:12AM

The site of LIB59 is UP.We will continue with the site of SEC59 in a few minutes.


Comment by OVH - Tuesday, 20 September 2011, 02:13AM

The site of SEC59 is UP.We will continue with the site of FAC59 in a moment.


Comment by OVH - Tuesday, 20 September 2011, 02:14AM

The site of FAC59 is UP.We will continue with the site of BOI59 in a few minutes.


Comment by OVH - Tuesday, 20 September 2011, 02:14AM

The site of BOI59 is UP. We will finish with the site BLE59.